# Distributed Hybrid Sketching for $\ell_{2}$-Embeddings

### v1

@@ V-B Local SRHTs with a Global Rademacher
Note on space complexity: Lastly, it is worth noting that our distributed approach is superior to any standard sketching approach in terms of space complexity. The largest matrix which is stored in one location at any given time throughout our procedure, is the initial global sketch $\widehat{\mathbf{A}}\in\mathbb{R}^{R\times d}$ (9), which is evidently smaller than $\mathbf{A}$. An artifact of this, is that our proposed approach is also beneficial in terms communication when we want to compute a sketch of distributed data, in the case when we would need to first aggregate the data in one server before applying the sketch.
VI Hybrid Sketching Experiments
In this section we consider several experiments to corroborate what we presented on hybrid sketching, for $\ell_{2}$-embeddings. All experiments were carried out on a single server, though the implementation of the the local sketches through $\widehat{\mathbf{S}}$ (8) can be carried out distributively and/or in parallel by utilizing $k$ server nodes.
VI-A Regression and $\ell_{2}$-s.e. Verification
Similar to the experiments of Sec. IV-D, for the following experiments we generated data matrices $\mathbf{A}\in\mathbb{R}^{N\times 40}$ following a $t$-distribution, where $N$ and $k$ is noted in the plots’ titles. We indicate the overall reduction factor by $\tau\coloneqq\rho\mu$, i.e. ${\tilde{R}}=\tau N$. We considered an initial reduction factor of $\mu$ varying between $0.5$ and $0.85$ in increments of $0.05$, and the second reduction was by a factor of $\rho=0.7$. In all the experiments of this subsection, it is clear that our error degrades gracefully, as $\tau$ decreases.
In Figure 6, we observe a clear benefit in terms of the empirically observed $\epsilon$ of (5) when $\mathbf{\Phi}$ is a Rademacher random sketch, for $k=8$. Furthermore, there seems to be a slight improvement when the local sketches are SRHTs, though when $\mathbf{\Phi}$ is an SRHT; the hybrid sketch is inferior to standard sketching approaches. This justifies our choice for $\mathbf{\Phi}$ and $\mathbf{S}_{\{k\}}$ from Sec. V-B.
Figure 6: $\ell_{2}$-s.e. error (5), for fixed $\rho=0.7$ and $\mu$ ranging between $0.7$ and $0.85$.
For the same experiment, in Figure 7 we considered $k=32$ and $N=3000$. Here, due to the fact that $N$ and $n$ are both significantly smaller, we observe a noticeable distinction when $\mathbf{S}_{\{k\}}$ are Rademacher compared to SRHT. In this case though, since $n$ is quite small, we do not have a significant benefit in terms of time complexity when it comes to local sketching. Here also, hybrid sketching outperforms standard sketching techniques. Similarly, we considered $N=8192$ with $k=32$, and analogous observations and conclusions were deduced. In this case, since $N$ was higher, we had better results. This is further justified by the fact that for higher $N$, the entries of $\widetilde{\mathbf{S}}$ provide a better approximation to Gaussian random variables.
Figure 7: $\ell_{2}$-s.e. error (5), for fixed $\rho=0.7$ and $\mu$ ranging between $0.7$ and $0.85$.
Figure 8: $\ell_{2}$-s.e. error (5), for fixed $\rho=0.7$ and $\mu$ ranging between $0.7$ and $0.85$.
VII Future Directions and Concluding Remarks
In the first portion of this work, we showed how we can obtain a global sketch of data distributed across a network in terms of a spectral guarantee, by performing local sketches and not aggregating until after the sketched matrices have taken place, satisfying desiderata $(\mathrm{a})$ and $(\mathrm{b})$ mentioned in the introduction. We also discussed the privacy aspect of this approach, as the local sketchings can provide security guarantees, meeting desideratum ($\mathrm{c}$). A potential future direction was mentioned in Sec. IV-A, in terms of improving the flattening of the leverage scores distributively. Another interesting avenue is investigating the ideas presented in this paper; when used for sparser/distributed Johnson-Lindenstrauss transforms, with partitions across the features of the data points.
Our distributive sketching approach was partly motivated by FL, where data is only available to local servers who may wish not to reveal their information. Our proposed approach can be utilized in first order federated algorithms to further accelerate them. It also permits exchange of local sketches for better approximations by correlating the sketches, while also providing a global summary of the data if desired. Moreover, by obtaining a compressed summary of the global dataset, one can administer global updates of mainstream federated approaches through distributed or centralized first and second order methods, while still meeting the objective of keeping the local data secret. Specific applications include PCA and low-rank recovery [69, 70], and subspace tracking [28]. To this extent, we also presented how our method can be utilized in distributed GD. Further investigation of these connections is worthwhile future work.
In the second part, we focused on hybrid sketching. Through hybrid sketching, we were able to satisfy all desiderata $(\mathrm{a})$-$(\mathrm{e})$. Though a simple idea, it has not been considered in previous works, other than few experiments carried out in [20]; with heuristic embeddings and reasoning, while lacking theoretical guarantees. As is evident, it is a powerful and beneficial technique; which should garner more attention, specially in distributed environments. We extensively discussed the combination we believe is ideal for hybrid sketching and the aforementioned desiderata, and provided numerical justification. This idea can be also be extended to sequentially applying multiple sketching matrices, for instance in training a neural network comprised of multiple training layers, which poses a bottleneck in contemporary machine learning. Motivated by a $k$-medians clustering algorithm [37], it would be interesting to see where else these ideas could be leveraged.
Appendix A Proofs of the main body
Proof.
[Lemma 1] Recall that for any $j\in\mathcal{I}_{i}$:
^$\hat{\ell}_{j}=\|\hat{\mathbf{U}}_{(j)}\|_{2}^{2}=\|\mathbf{e}_{i}^{\top}\hat{\mathbf{U}}\|_{2}^{2}=\mathbf{e}_{j}^{\top}\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top}\mathbf{e}_{j}\ .$
It then follows that
$=\frac{\mathscr{L}_{i}}{n}\ .$
Proof.
[Proposition 1] We know that ^$\hat{\ell}_{j}\in[0,\mathscr{L}_{i}]$, and the normalized scores are ¯$\bar{\ell}_{j}=\hat{\ell}_{j}/d$ for each $j\in\mathcal{I}_{i}$. By Lemma 1, it follows that
Proof.
for which $\Pr[B_{ji}]=1/k$ and $\mathbb{E}[B_{ji}]=1/k$, as $B_{ji}$ follows a Bernoulli distribution. Further define $Y^{i}\coloneqq\sum_{j=1}^{R}B_{ji}$; which follows a Binomial distribution, hence
Proof.
[Theorem 2] For each $i\in\mathbb{N}_{k}$ and every $j\in\mathcal{I}_{i}$, by Proposition 1 we can assume that ¯$\bar{\ell}_{j}\approx\mathscr{L}_{i}/(nd)$ (w.h.p.), where ¯$\bar{\ell}_{j}$ is the sampling probability $\tilde{\pi}_{j}$ that row $j$ is sampled at each independent trial. Hence, Algorithm 1 is now performing approximate leverage score sampling with distribution $\tilde{\mathcal{U}}_{N}=\bigcup_{\iota=1}^{k}\left(\bigcup_{j\in\mathcal{I}_{\iota}}\big{\{}\mathscr{L}_{\iota}/(nd)\big{\}}\right)$. The misestimation factor for sampling according to $\mathcal{U}_{N}$ instead of $\tilde{\mathcal{U}}_{N}$, is
By Proposition 2, the resulting $\hat{\mathbf{\Omega}}$ through the local sampling matrices $\mathbf{\Omega}_{\{k\}}$, is close to a uniform sampling matrix (w.r.) of $R$ out of $N$ elements. By applying the above conclusions to Theorem 1, the proof is complete. ∎
Proof.
Proof.
Proof.
Proof.
We denote the resulting sketching matrix by $\widehat{\mathbf{S}}_{[t]}$. Since in $\tilde{\mathbf{g}}^{[t]}$ the only source of randomness comes from $\widehat{\mathbf{S}}_{[t]}$, we have $\mathbb{E}\big{[}\tilde{\mathbf{g}}^{[t]}\big{]}=2\mathbf{A}^{\top}\mathbb{E}\big{[}\widehat{\mathbf{S}}_{[t]}^{\top}\widehat{\mathbf{S}}_{[t]}\big{]}(\mathbf{A}\mathbf{x}^{[t]}-\mathbf{b})$. By modifying the proof of [50, Lemma 1], we can show that $\mathbb{E}\big{[}\widehat{\mathbf{S}}_{[t]}^{\top}\widehat{\mathbf{S}}_{[t]}\big{]}=\mathbf{I}_{N}$. It suffices to show that $\mathbb{E}\big{[}\hat{\mathbf{\Omega}}_{[t]}^{\top}\hat{\mathbf{\Omega}}_{[t]}\big{]}=\mathbf{I}_{N}$. Let $\mathcal{J}_{i}$ denote the index set of $\hat{\mathbf{\Omega}}_{[t]}$’s rows corresponding to $\mathbf{\Omega}_{i}^{[t]}$, i.e. $\mathcal{J}_{i}\subset\mathbb{N}_{R}$; $\bigsqcup_{\iota=1}^{k}\mathcal{J}_{\iota}=\mathbb{N}_{R}$; and #$\#\mathcal{J}_{i}=r$ for each $i\in\mathbb{N}_{k}$. Then, $\hat{\mathbf{\Omega}}_{[t]}\big{|}_{(\mathcal{J}_{i})}$ is $\mathbf{\Omega}_{i}^{[t]}$ presented in the submatrix corresponding to rows indexed by $\mathcal{J}_{i}$, and zeros everywhere else. We then have
$=\sum_{i=1}^{k}\mathbf{I}_{N}\big{|}_{(\mathcal{J}_{i})}=\mathbf{I}_{N}$
and the proof is complete. ∎
Proof.
For now, let $(\epsilon_{1},\delta_{1})=(\epsilon_{2},\delta_{2})=({\tilde{\epsilon}},\tilde{\delta})$ denote the parameters of the corresponding sketching matrices. By the assumption on $\mathbf{\Pi}_{1}$, for any $\mathbf{x}$ we have $\Pr\left[\|\mathbf{\Pi}_{1}\mathbf{x}\|>_{\epsilon_{1}}\|\mathbf{x}\|\right]<\delta_{1}$. Further define the set of all $\ell_{2}$-s.e. of $\mathbf{U}$ with distortion parameter $\epsilon$:
$\mathcal{E}_{\mathbf{U}}^{\epsilon}\coloneqq\left\{S_{i}\ :\ S_{i}\text{ is an}\epsilon\ \ell_{2}\text{-s.e. of }\mathbf{U}\right\}.$
$\overset{\diamondsuit}{<}\delta_{2}\cdot 1+\delta^{\prime}$
$\overset{\heartsuit}{<}\delta_{1}+\delta_{2}\ .$
In $\natural$, we used the fact that the conditional probability in the second summand is always less than 1, and the fact that for an embedding with a smaller distortion; we have a higher error probability, i.e. $\delta^{\prime}<\delta_{2}$ since $\epsilon^{\prime}>{\tilde{\epsilon}}=\epsilon_{2}$. In $\flat$ we use the fact that the conditional probability in the first summand corresponds to the failure event of $\mathbf{\Pi}_{2}$ not being an $\epsilon_{2}$ $\ell_{2}$-s.e., which occurs with probability less than $\delta_{2}$. In $\sharp$ we concisely represent the two summations, which are equal to the depicted events. In turn, we substitute their upper bounds which are respectively 1 and $\delta^{\prime}$. Lastly, in $\heartsuit$ we use the fact that $\delta_{1}>\delta^{\prime}$. By our choice of $\tilde{\delta}$, we know that $\delta_{1}+\delta_{2}=\delta$.
Proof.
[Theorem 3] By Theorem 2 and Corollary 1, we know that our distributed local sketching approach through Algorithm 1 can be used to obtain an $({\tilde{\epsilon}},\tilde{\delta})$ $\ell_{2}$-s.e. sketch $\widehat{\mathbf{A}}=\widehat{\mathbf{S}}\mathbf{A}$ of $\mathbf{A}$. Then, by applying another $({\tilde{\epsilon}},\tilde{\delta})$ $\ell_{2}$-s.e. sketching matrix $\mathbf{\Phi}$ on $\widehat{\mathbf{A}}$, according to Proposition 6, we obtain an $(\epsilon,\delta)$ $\ell_{2}$-s.e. sketch $\widetilde{\mathbf{A}}=\mathbf{\Phi}\cdot\widehat{\mathbf{A}}=(\mathbf{\Phi}\widetilde{\mathbf{S}})\cdot\mathbf{A}$ of $\mathbf{A}$. This completes the proof. ∎
Appendix B Gradient Descent Through Distributed Local Sketching
Here we provide the details to how our proposed scheme on distributed local sketching can be utilized to solve a linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ in which the data is itself distributed through GD, which was discussed in Sec. IV-C.
@@ Appendix B Gradient Descent Through Distributed Local Sketching
First, recall that the gradient of $L_{ls}(\mathbf{A},\mathbf{b};\mathbf{x})\coloneqq\|\mathbf{A}\mathbf{x}-\mathbf{b}\|_{2}^{2}$ defined in (1) for a given $\mathbf{x}^{[t]}$ at iteration $t$ of GD, is
@@
To carry out GD distributively, through iterative sketching, the local servers will perform a distinct $\big{\{}\mathbf{\Omega}^{[t]}_{i}\big{\}}_{i=1}^{k}$ at each iteration $t$, to get a different local sketches
@@
of their local dataset. Note that at each iteration, the local label vectors $\mathbf{b}_{\{k\}}$ are sketched accordingly, with the same corresponding sketching matrices. The pseudocode of this approach is summarized in Algorithm 3.
@@
The expected regret of the least squares objective resulting from distributed GD with iterative sketching; according to Algorithm 3, with a diminishing step-size $\eta_{s}$, converges to zero at a rate of $\mathcal{O}(1/\sqrt{s}+R/s)$, i.e.
@@
       1.0)1.0)1.0 ) The coordinator delivers $r$ and ${\hat{\mathbf{x}}}^{[0]}$ to the $k$ servers
@@
       1.1)\ $)1$ Generate a random $\mathbf{P}_{i}\in\mathbb{R}^{n\times n}$
       1.2)\ 1.2)1.2 ) Locally transform the data [^4] To simplify the pseudocode presentation, in contrast to the above description, we apply the rescaling simultaneously with $\mathbf{P}_{\{k\}}$.:
@@
             2.1)\ 2.1)2.1 ) Uniformly sample $r$ rows from $\acute{\mathbf{A}}_{i}$, through $\mathbf{\Omega}_{i}^{[t]}$:
@@
             2.2)\ $)2$ Determine the sketched partial gradient:
@@
             3.1)\ 3.1)3.1 ) Determines $\eta_{t}$ and the global sketched gradient:
@@
             3.2)\ 3.2)3.2 ) GD Update: ${\hat{\mathbf{x}}}^{[t+1]}\leftarrow{\hat{\mathbf{x}}}^{[t]}-\eta_{t}\cdot\hat{g}^{[t]}$
             3.3)\ $)3$ Delivers ${\hat{\mathbf{x}}}^{[t]}$ to the $k$ servers

### v2

@@ V-B Local SRHTs with a Global Rademacher
Space requirement: It is worth noting that our distributed approach is superior to any standard sketching approach in terms of space requirement. The largest matrix which is stored in one location at any given time throughout our procedure, is the initial global sketch $\widehat{\mathbf{A}}\in\mathbb{R}^{R\times d}$ (9), which is evidently smaller than $\mathbf{A}$. An artifact of this, is that our proposed approach is also beneficial in terms communication when we want to compute a sketch of distributed data, in the case when we would need to first aggregate the data in one server before applying the sketch.
Network constraints: Lastly, we outline the parameters selection when we have network imposed constraints, such as the number of servers, communication and storage budgets. We can assume that $k$ is fixed and all servers are homogeneous, with storage capacity of at least $n$ data points of length $d$. If we now assume a communication budget of $C_{1}$ symbols from each server to the coordinator, in order to maintain as much information as possible from the local sketches and utilize the allowable network communication, we set $\mu=C_{1}/(nd)$. If the coordinator has a storage budget of $C_{2}$ symbols, we have to set $\mu=\min\left\{C_{1}/(nd),C_{2}/(knd)\right\}$. Upon receiving the local sketches, the coordinator has a sketch $\widehat{\mathbf{A}}\in\mathbb{R}^{R\times d}$ according to (9) where $R=\lfloor\mu N\rfloor{}$. Then, if the final sketch $\widetilde{\mathbf{A}}$ has a further constraint of not exceeding $C_{3}$ symbols, we will require that the second sketching matrix $\mathbf{\Phi}$ is of size ${\tilde{R}}\times R$, where $\rho=C_{3}/(Rd)$ and ${\tilde{R}}=\lfloor\rho R\rfloor$. It is important to note that the choice of sketching method and associated parameters directly influences the accuracy guarantees $(\epsilon,\delta)$ of $\widehat{\mathbf{A}}$. For this reason, we seek to retain as many data points as possible at each step.
VI Distributed Gradient Descent and Security
VI-A Implications to Distributed Gradient Descent
An implication of [22, Thm. 3], is that as long as the $\ell_{2}$-s.e. property is satisfied by the iterative sketching [^2] By iterative sketching, we mean that at each iteration of the algorithm, we apply a new sketching matrix. (or sampling) technique which takes place, we converge in expectation to the optimal solution (1) at a rate inversely proportional to the number of iterations. To establish that our sketching technique (Fig. 1) can be utilized to obtain the same convergence guarantee, it suffices to show that if every server performs a different sampling $\mathbf{\Omega}_{i}^{[t]}$ at each iteration $t$, i.e. a new global underlying sketching matrix $\widehat{\mathbf{S}}^{[t]}$ is applied at each iteration, the overall gradient is an unbiased estimate of the gradient $\nabla_{\mathbf{x}}L_{ls}(\mathbf{A},\mathbf{b};\mathbf{x}^{[t]})$, which we show this in Proposition 5.
Proposition 5.
Corollary 2 ([22, Thm. 3]).
Assume a bounded gradient on the least squares objective resulting from distributed GD with iterative sketching according to Algorithm 1, with step-size $\eta_{t}=1/(\alpha t)$ for a fixed $\alpha>0$. Then, after $T$ iterations:
It is worth mentioning that even though we directly invoke [22, Thm. 3] to give a convergence guarantee when the central server aggregates the gradients and performs distributed gradient descent, the formulation here is vastly different to that of previous works. In those works, the projections are performed on the aggregated data, which is the main contribution of this section, i.e. we showed that you do not need to aggregate the data to get a global embedding. Furthermore, this application to distributed GD can be viewed as an extension of the prior works, where now the data is meant to be kept private from the coordinator rather than the servers. Additionally, here we require less overall storage, as the coordinator only needs to keep track of the updated estimates $\tilde{\mathbf{g}}^{[t]}$ and ${\hat{\mathbf{x}}}^{[t+1]}={\hat{\mathbf{x}}}^{[t]}-\eta_{t}\cdot{\hat{\mathbf{x}}}^{[t]}$, for $\eta_{t}$ an appropriate step-size.
@@ VI-A Implications to Distributed Gradient Descent
Recall that the gradient of $L_{ls}(\mathbf{A},\mathbf{b};\mathbf{x})\coloneqq\|\mathbf{A}\mathbf{x}-\mathbf{b}\|_{2}^{2}$ defined in (1) for a given $\mathbf{x}^{[t]}$ at iteration $t$ of GD, is
@@
To carry out GD distributively, through iterative sketching, the local servers will perform a distinct $\big\{\mathbf{\Omega}^{[t]}_{i}\big\}_{i=1}^{k}$ at each iteration $t$, to get different local sketches
@@
of their local dataset. Note that at each iteration, the local label vectors $\mathbf{b}_{\{k\}}$ are sketched accordingly, with the same corresponding sketching matrices. The pseudocode of this approach is summarized in Algorithm 3. To simplify the pseudocode presentation, in contrast to the above description, we apply the rescaling simultaneously with $\mathbf{P}_{\{k\}}$.
@@
The expected regret of the least squares objective resulting from distributed GD with iterative sketching according to Algorithm 3, with a diminishing step-size $\eta_{s}$, converges to zero at a rate of $\mathcal{O}(1/\sqrt{s}+R/s)$, i.e.
@@
 $1.0)$ The coordinator delivers $r$ and ${\hat{\mathbf{x}}}^{[0]}$ to the $k$ servers
@@
  1.1)\ 1.1) Generate a random $\mathbf{P}_{i}\in\mathbb{R}^{n\times n}$
  1.2)\ 1.2) Locally transform the data:
@@
     2.1)\ 2.1) Uniformly sample $r$ rows from $\acute{\mathbf{A}}_{i}$, through $\mathbf{\Omega}_{i}^{[t]}$:
@@
     2.2)\ 2.2) Determine the sketched partial gradient:
@@
     3.1)\ 3.1) Determines $\eta_{t}$ and the global sketched gradient:
@@
     3.2)\ 3.2) GD Update: ${\hat{\mathbf{x}}}^{[t+1]}\leftarrow{\hat{\mathbf{x}}}^{[t]}-\eta_{t}\cdot\hat{g}^{[t]}$
     3.3)\ 3.3) Delivers ${\hat{\mathbf{x}}}^{[t]}$ to the $k$ servers